# Fluctuations and the limit of predictability in protein evolution

### v2

@@ Abstract
Protein evolution involves mutations occurring across a wide range of time scales. In analogy with other disordered systems, this dynamical heterogeneity suggests strong correlations between mutations happening at distinct sites and times. To quantify these correlations, we examine the role of various fluctuation sources in protein evolution, simulated using a data-driven epistatic landscape. By applying spatio-temporal correlation functions inspired by statistical physics, we disentangle fluctuations originating from the ancestral protein sequence from those driven by stochastic mutations along independent evolutionary paths. Our analysis shows that, in diverse protein families, fluctuations from the ancestral sequence predominate at shorter time scales. This allows us to identify a time scale over which ancestral sequence information persists, enabling its reconstruction. We link this persistence to the strength of epistatic interactions: ancestral sequences with stronger epistatic signatures impact evolutionary trajectories over extended periods. At longer time scales, however, ancestral influence fades as epistatically constrained sites evolve collectively. To confirm this idea, we apply a standard ancestral sequence reconstruction algorithm and verify that the time-dependent recovery error is influenced by the properties of the ancestor itself.
@@ I Introduction
Proteins play a key role in many biological processes that are essential for life. At the same time, they display a huge evolutionary flexibility, in that a large diversity of protein sequences can fold into the same structure and perform the same biological function. Such proteins are called ‘homologous’ and grouped into the same ‘protein family’ Blum et al. (2024); Paysan-Lafosse et al. (2024); Consortium (2022); Burley et al. (2022). During evolution, a single ancestral protein belonging to a given family can diversify its amino acid sequence and, given enough time, explore a portion of the ‘neutral space’ of equivalently fit protein sequences Kimura (1985). Nevertheless, a random mutation in the amino acid sequence has a high probability of negatively impacting the structure and functionality of the protein Fowler and Fields (2014); Notin et al. (2024). Such kind of deleterious mutations are eliminated by natural selection, by which certain protein variants are favored over others due to their functional advantages: neutral (or even beneficial) mutations that maintain (or enhance) protein structure, stability, or function are more likely to persist in a population Hartl et al. (1997).
@@
The picture gets more complex as one takes into account the concept of epistasis De Visser and Krug (2014); Harms and Thornton (2013); Starr and Thornton (2016); Domingo et al. (2019); Johnson et al. (2023); Lunzer et al. (2010); Biswas et al. (2019); Poelwijk et al. (2019); Chen et al. (2023); Poelwijk et al. (2016); Reddy and Desai (2021); Otwinowski (2018): the effect of a mutation on the sequence fitness changes with the ‘background’ in which the mutation takes place, i.e. the amino acids that are present in the other sites of the protein. As a consequence, a mutation that would be deleterious in a particular background sequence can be instead beneficial in another (also called sign epistasis), allowing evolution to explore different pathways.
@@
Understanding and characterizing the impact of epistasis in evolution requires careful experiments and modeling. In particular, recent developments increased substantially the sequence divergence that can be reached by laboratory evolution experiments Fantini et al. (2020); Stiffler et al. (2020); Erdoğan et al. (2023); Park et al. (2022); Rix et al. (2023), such that many data are becoming available and many more are expected to come soon. Yet, the sequence diversity of natural evolution still remains out of reach of such experiments, thus leaving an unexplored gap in evolutionary time scales. In order to fill this gap, one can simulate the evolution of protein sequences in silico de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2024); Biswas et al. (2024); Bari et al. (2024), relying on the data-driven approach that goes under the name of Direct Coupling Analysis (DCA) Morcos et al. (2011); Cocco et al. (2018), in which a fitness landscape (analogous to an energy function in the statistical physics vocabulary) is inferred starting from a Multiple Sequence Alignment (MSA) of natural homologs constituting a given protein family Ferguson et al. (2013); Figliuzzi et al. (2016); Levy et al. (2017); Couce et al. (2017); Vigué and Tenaillon (2023). The energy function that results from this inference procedure can then be used to assign a probability to each sequence. The resulting landscape is explored by means of a biologically motivated Monte Carlo Markov-Chain (MCMC) algorithm, providing in silico evolutionary trajectories that quantitatively mimic experimental results de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2022); Biswas et al. (2024); Bari et al. (2024).
@@
Within this framework, it has been recently shown Bari et al. (2024) that different sites evolve with widely different time scales, which also depend on the background sequence, due to epistatic interactions. This heterogeneity of local time scales is strongly reminiscent of similar phenomena observed in disordered physical systems Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Berthier and Jack (2007); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018). There, it has been shown that a large heterogeneity in time scales implies the presence of strong and highly collective space-time correlations between sites that slow down the dynamical evolution. Furthermore, it has been shown that the structural disorder in the initial configuration partially encodes future correlations, and can be used (possibly by machine learning tools) to predict the future dynamics Widmer-Cooper et al. (2008); Schoenholz et al. (2016); Bapst et al. (2020); Jung et al. (2024), see Jung et al. (2023) for a recent review. These are the main observations that motivated this work.
@@
In this paper, we measure space-time correlations to explore how the heterogeneity of evolutionary time scales relates to the epistatic interactions within the ancestral sequence. We follow the dynamics of the Hamming distance, i.e., the number of accepted mutations with respect to the ancestral sequence. We characterize how this quantity fluctuates (i) between different evolutionary trajectories originating from the same ancestor, and (ii) between different ancestors. We observe a strong dependence of the dynamics on the ancestral sequence for short enough times, over which the first source of fluctuations is found to be subdominant with respect to the second. On the other hand, at long times the stochasticity of evolutionary trajectories takes over, and any memory of the ancestor is lost. This has important consequences, as it allows us to properly quantify the time scale over which it should be possible to reconstruct the ancestral sequence of a certain set of evolutionary trajectories, and how this time scale depends on the epistatic interactions in the ancestor itself. In fact, we find that the amount of epistatically constrained sites in the ancestral sequence determines this time scale: more epistatic sequences leave their trace on evolutionary dynamics for longer times. We then measure the correlations between the evolution of all pairs of sites at the time scale at which such correlations are stronger, finding a different pattern for each ancestral sequence, that is then reflected in the evolutionary dynamics. Finally, we show that the magnitude of epistatic interactions, which controls the stochasticity of the evolutionary trajectories, is also related to the linear response of the evolutionary dynamics to a change in selective pressure (controlled by the temperature in our Monte Carlo simulations). Hence, we show and quantify how more epistatically constrained ancestors lead to a more complex evolutionary dynamics over longer time scales, which is also more sensitive to perturbations of the environment.
II Methods
II.1 Modeling evolution in silico
@@
Following Bari et al. (2024); de la Paz et al. (2020); Bisardi et al. (2022), we consider a fixed initial sequence $A_{0}=(a^{0}_{1},\dots,a^{0}_{L})$ as the ancestor, and we let many trajectories evolve from it in parallel, hence restricting ourselves to a star phylogeny describing an ensemble of independent evolutionary trajectories of common initialization. Our goal is to characterize the statistical properties of the specific ancestral sequence and the sequence space accessible from it in a given evolutionary time. The sequence evolution is simulated by a Monte Carlo dynamics. In this work, we use the Metropolis algorithm acting on amino acids for simplicity, and we verified that more refined sampling strategies taking into account amino-acid accessibility via the genetic code, insertions and deletions Bari et al. (2024); de la Paz et al. (2020); Bisardi et al. (2022) produce qualitatively the same results (see SM). As it was previously shown Russ et al. (2020), at large times the generated sequences accurately reproduces many statistical features of the natural sequences used for the training, and have the same probability of being biologically functional in a given experimental platform; hence, the model is generative. Here, we are concerned with what happens at short and intermediate times, where the influence of the ancestral sequence is still important.
@@
II.2 Measures of epistasis
By using as input data the MSA of a given protein family, and inferring the fitness landscape parameters via the DCA model, the authors of Vigué et al. (2022); Bari et al. (2024) have been able to roughly classify the sites of any specific protein sequence belonging to the family in three categories: conserved, mutable, and epistatically constrained. In order to do so, they defined two site-mutability metrics.
@@ II.2 Measures of epistasis
In terms of these quantities, a site $i$ in a given background sequence $A$ can thus be classified as follows.
@@
Furthermore, and most importantly for the present work, Ref. Bari et al. (2024) has considered a given background $A_{0}$ as the ancestral sequence, and starting from it has performed many parallel evolutions in silico, looking at how each site diversifies in the library of mutants obtained after a certain evolutionary time. It was found that mutable sites evolve very rapidly and quickly reach their asymptotic mutability, i.e. the CIE. Conserved sites remain so at any time during evolution, hence do not display interesting dynamics. Epistatically constrained sites, instead, are very conserved at short evolutionary times, due to the epistatic constraints in the background of the ancestor, which cause a small CDE. But, as soon as the background sequence mutates enough, they can tolerate more mutations, asymptotically reaching the large CIE that they display in the natural alignment. Their mutation, however, is contingent on several mutations happening in the rest of the sequence, which often require a rather long evolutionary time to take place Bari et al. (2024).
@@
Figure 1: Schematic representation of (a) a MSA in which each site is colored based on its mutability as measured by the $\textrm{CDE}_{i}^{A}$ and (b) the protocol of our evolutionary dynamics with the relevant averages. (c) Depending on the epistatic constraints acting on the ancestral sequence and for equal evolutionary times, we observe a different level of diversity of the evolved sequences (colored), corresponding to a different exploration of the functional sequence space (grey).
II.3 Dynamical fluctuations
@@
Next, we consider a star phylogeny of independent MCMC evolutionary trajectories, all starting from the same ancestor and evolving in parallel (Fig. 1b). Hence, for each ancestor, we construct, as a function of evolutionary time, a MSA of descendants that mimic those obtained through in vitro evolution experiments. As a measure of diversity, we focus on the evolution of the Hamming distance, between the ancestor $A_{0}$ and the MSA of evolving sequences at time $t$, $A_{t}=(a^{t}_{1},\cdots,a^{t}_{L})$, defined as
@@
To make this observation quantitative, we characterize the statistical properties of the resulting MSAs, and how they depend on the ancestor, by introducing two distinct averages and corresponding fluctuations, inspired from the disordered systems literature Franz et al. (2011, 2013); Seoane and Zamponi (2018); Berthier and Jack (2007) and illustrated in Fig. 1b. Recall that the number of accepted mutations $H^{A_{0}}(t)$ at evolutionary time $t$ for fixed ancestor $A_{0}$ is a random variabile, whose realizations depend on the stochasticity of the evolutionary trajectory.
@@
is the average of the ancestor-dependent $\chi^{A_{0}}_{\rm dyn}(t)$ over $A_{0}$. Using these quantities, which are called ‘dynamical susceptibilities’ in the physics of disordered systems, we can thus quantify the relative importance of the ancestor and the evolutionary noise in determining the diversity of the resulting MSAs at a fixed evolutionary time. In the SM we also study some alternative definitions of susceptibilities, based for example on the Hamming distance between two chains starting from the same ancestor. However, because our MCMC dynamics is time-reversible, these two quantities are related.
III Results
III.1 Mutational dynamics and its fluctuations

### v3

@@ Abstract
Protein evolution involves mutations occurring across a wide range of time scales. In analogy with disordered systems in statistical physics, this dynamical heterogeneity suggests strong correlations between mutations happening at distinct sites and times. To quantify these correlations, we examine the role of various fluctuation sources in protein evolution, simulated using a data-driven energy landscape as a proxy for protein fitness. By applying spatio-temporal correlation functions developed in the context of disordered physical systems, we disentangle fluctuations originating from the initial condition, i.e. the ancestral sequence from which the evolutionary process originated, from those driven by stochastic mutations along independent evolutionary paths. Our analysis shows that, in diverse protein families, fluctuations from the ancestral sequence predominate at shorter time scales. This allows us to identify a time scale over which ancestral sequence information persists, enabling its reconstruction. We link this persistence to the strength of epistatic interactions: ancestral sequences with stronger epistatic signatures impact evolutionary trajectories over extended periods. At longer time scales, however, ancestral influence fades as epistatically constrained sites evolve collectively. To confirm this idea, we apply a standard ancestral sequence reconstruction algorithm and verify that the time-dependent recovery error is influenced by the properties of the ancestor itself. Overall, our results reveal that the properties of ancestral sequences—particularly their epistatic constraints—influence the initial evolutionary dynamics and the performance of standard ancestral sequence reconstruction algorithms.
@@ I Introduction
Proteins play a key role in many biological processes that are essential for life. At the same time, they display a huge evolutionary flexibility, in that a large diversity of protein sequences can fold into the same structure and perform the same biological function. Such proteins are called ‘homologous’ and grouped into the same ‘protein family’ Blum et al. (2024); Paysan-Lafosse et al. (2024); Consortium (2022); Burley et al. (2022). During evolution, a single ancestral protein belonging to a given family can diversify its amino acid sequence and, given enough time, explore a portion of the ‘neutral space’ of equivalently fit protein sequences Kimura (1983). Nevertheless, a random mutation in the amino acid sequence has a high probability of negatively impacting the structure and functionality of the protein Fowler and Fields (2014); Sarkisyan et al. (2016); Notin et al. (2024). Such kind of deleterious mutations are eliminated by natural selection, by which certain protein variants are favored over others due to their functional advantages: neutral (or even beneficial) mutations that maintain (or enhance) protein structure, stability, or function are more likely to persist in a population Hartl et al. (1997).
@@
The picture gets more complex as one takes into account the concept of epistasis Harms and Thornton (2013); Weinreich et al. (2013); De Visser and Krug (2014); Starr and Thornton (2016); Poelwijk et al. (2016); Cocco et al. (2018); Domingo et al. (2019); Johnson et al. (2023); Buda et al. (2023): the effect of a mutation on the sequence fitness changes with the ‘background’ in which the mutation takes place, i.e. the amino acids that are present in the other sites of the protein. As a consequence, a mutation that would be deleterious in a particular background sequence can be instead beneficial in another (also called sign epistasis), allowing evolution to explore different pathways.
@@
Understanding and characterizing the impact of epistasis in evolution requires careful experiments and modeling. In particular, recent developments have substantially increased the sequence divergence that can be reached by laboratory evolution experiments Fantini et al. (2020); Stiffler et al. (2020); Erdoğan et al. (2024); Park et al. (2022); Rix et al. (2024). A large amount of data is therefore now becoming available with more expected to come soon. Yet, the sequence diversity of natural evolution still remains out of reach of such experiments, thus leaving an unexplored gap in evolutionary time scales. In order to fill this gap, one can simulate the evolution of protein sequences in silico de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2022, 2024); Biswas et al. (2024); Bari et al. (2024), relying on the data-driven approach that goes under the name of Direct Coupling Analysis (DCA) Morcos et al. (2011); Cocco et al. (2018). DCA infers a fitness landscape (analogous to an energy function in the statistical physics vocabulary) starting from a Multiple Sequence Alignment (MSA) of natural homologs constituting a given protein family Ferguson et al. (2013); Figliuzzi et al. (2016); Levy et al. (2017); Couce et al. (2017); Vigué and Tenaillon (2023). The energy function that results from this inference procedure can then be used to assign a probability to each sequence. The resulting landscape is explored by means of a biologically motivated Monte Carlo Markov-Chain (MCMC) algorithm, providing in silico evolutionary trajectories that quantitatively mimic experimental results Biswas et al. (2019); de la Paz et al. (2020); Bisardi et al. (2022); Alvarez et al. (2024, 2022); Biswas et al. (2024); Bari et al. (2024).
@@
Within this framework, it has been recently shown Bari et al. (2024) that different sites evolve with widely different time scales, which also depend on the background sequence, due to epistatic interactions. More precisely, sites that are more epistatically constrained need much longer times (or number of generations) to evolve (i.e. accumulate mutations) compared to less constrained sites Bari et al. (2024). Furthermore, whether a site is epistatically constrained or not depends on the rest of the sequence (the background) Lyons et al. (2020); Vigué et al. (2022); Chen et al. (2024), which adds sequence-to-sequence heterogeneity on top of the site-to-site heterogeneity. This “heterogeneity of time scales” is strongly reminiscent of similar phenomena observed in disordered physical systems Kirkpatrick and Thirumalai (1988); Franz et al. (1999); Franz and Parisi (2000); Bouchaud and Biroli (2005); Berthier et al. (2005); Berthier and Jack (2007); Franz et al. (2011); Berthier et al. (2011); Franz et al. (2013); Seoane and Zamponi (2018); Folena et al. (2022). There, it has been shown that a large heterogeneity in time scales implies the presence of strong and highly collective space-time correlations between sites that slow down the dynamical evolution. Furthermore, it has been shown that the structural disorder in the initial configuration partially encodes future correlations, and can be used (possibly by machine learning tools) to predict the future dynamics Widmer-Cooper et al. (2008); Schoenholz et al. (2016); Bapst et al. (2020); Jung et al. (2024), see Jung et al. (2025) for a recent review. These are the main observations that motivated this work.
@@
In this paper, we measure space-time correlations to explore how the heterogeneity of evolutionary time scales relates to the epistatic interactions within the ancestral sequence. We follow the dynamics of the Hamming distance, i.e., the number of accepted mutations with respect to the ancestral sequence. We characterize how this quantity fluctuates (i) between different evolutionary trajectories originating from the same ancestor, and (ii) between different ancestors. We observe a strong dependence of the dynamics on the ancestral sequence for short enough times, over which the first source of fluctuations is found to be subdominant with respect to the second. On the other hand, at long times the stochasticity of evolutionary trajectories takes over, and any memory of the ancestor is lost. This has important consequences, as it allows us to properly quantify the time scale over which it should be possible to reconstruct the ancestral sequence of a certain set of evolutionary trajectories, and how this time scale depends on the epistatic interactions in the ancestor itself. In fact, we find that the amount of epistatically constrained sites in the ancestral sequence determines this time scale: more epistatic sequences leave their trace on evolutionary dynamics for longer times. We then measure the correlations between the evolution of all pairs of sites at the time scale at which such correlations are stronger, finding a different pattern for each ancestral sequence, which is then reflected in the evolutionary dynamics. Finally, we show that the magnitude of epistatic interactions, which controls the stochasticity of the evolutionary trajectories, is also related to the linear response of the evolutionary dynamics to a change in selective pressure (controlled by the temperature in our Monte Carlo simulations). Hence, we show and quantify how more epistatically constrained ancestors lead to a more complex evolutionary dynamics over longer time scales, which is also more sensitive to perturbations of the environment.
II Epistasis in biological data
Epistasis plays a central role in shaping protein evolution, influencing both the structure of fitness landscapes and the fate of evolutionary trajectories, and as such it has been the subject of theoretical, computational and experimental studies, see Refs. Harms and Thornton (2013); Weinreich et al. (2013); De Visser and Krug (2014); Starr and Thornton (2016); Poelwijk et al. (2016); Cocco et al. (2018); Domingo et al. (2019); Johnson et al. (2023); Buda et al. (2023) for a few recent reviews. It has been shown that some epistatic effects can be accounted for by a global non-linearity of the phenotype-fitness relation Otwinowski (2018); Otwinowski et al. (2018); Reddy and Desai (2021); Schulte et al. (2025), but epistasis is also due to a network of more specific pairwise and higher-order interactions. While some of these interactions are sparse Sailer and Harms (2017a, b); Domingo et al. (2018); Poelwijk et al. (2019); Ballal et al. (2020); Phillips et al. (2021); Buda et al. (2023); Miton et al. (2021), multiple studies suggest that the most relevant effects tend to emerge collectively, from the accumulation of numerous weak interactions between a single site and multiple other residues across the protein sequence Lunzer et al. (2010); Miton and Tokuriki (2016); Figliuzzi et al. (2016); Rivoire et al. (2016); Poelwijk et al. (2016); Cocco et al. (2018); Starr et al. (2018); Vigué et al. (2022); Chen et al. (2024). This form of distributed epistasis suggests that the evolutionary constraints on a given mutation are shaped by the broader sequence context, rather than by a few dominant interactions, highlighting the complexity of protein fitness landscapes.
Experiments have provided valuable insights into how mutations interact, but each methodology comes with inherent limitations. While some studies focus on measuring epistasis in a restricted mutational space, others provide broader datasets but lack information on evolutionary dynamics. Here, we summarize the main categories of experimental data available and discuss their relevance to our theoretical framework.
  • Combinatorial mutagenesis Poelwijk et al. (2019); Phillips et al. (2021); Bakerlee et al. (2022); Buda et al. (2023); Papkou et al. (2023); Somermeyer et al. (2022); Schulz et al. (2025) experiments involving a small number of residues have demonstrated epistatic effects. However, these studies typically do not allow the accumulation of sufficient mutational effects to observe the large-scale evolutionary patterns we investigate in this work.
  • Deep mutational scanning (DMS) experiments across homologous wild-type proteins Olson et al. (2014); Park et al. (2022); Chen et al. (2024); Romanowicz et al. (2025), reveal epistatic interactions and can be used to explore the concept of site variability within a small region of the landscape. However, these experiments lack direct observations of evolutionary dynamics over multiple generations.
  • Experimental studies tracking evolutionary dynamics in vitro are available Fantini et al. (2020); Stiffler et al. (2020); Erdoğan et al. (2024); Rix et al. (2024), but only a few cases, such as TEM-1 Fantini et al. (2020) and PSE-1 Stiffler et al. (2020) $\beta$-lactamases, involve multiple wild-type homologs belonging to the same protein family, whose evolutionary dynamics can be directly compared. Moreover, these experiments do not extend far enough in sequence space to capture the long-term epistatic effects central to our study.
Given these limitations, we anticipate that future advancements in experimental techniques will provide richer datasets capable of testing the predictions outlined in our study. In the meantime, our theoretical framework serves as a guide for interpreting existing data and shaping expectations for future experimental work.
III Methods
III.1 Modeling evolution in silico
@@
Following Bari et al. (2024); de la Paz et al. (2020); Bisardi et al. (2022), we consider a fixed initial sequence $A_{0}=(a^{0}_{1},\dots,a^{0}_{L})$ as the ancestor, and we let many trajectories evolve from it in parallel, hence restricting ourselves to a star phylogeny describing an ensemble of independent evolutionary trajectories of common initialization. Our goal is to characterize the statistical properties of the specific ancestral sequence and the sequence space accessible from it in a given evolutionary time. The sequence evolution is simulated by Monte Carlo dynamics. In this work, we use the Metropolis algorithm acting on amino acids for simplicity, and we verified that more refined sampling strategies taking into account amino-acid accessibility via the genetic code, insertions and deletions Bari et al. (2024); de la Paz et al. (2020); Bisardi et al. (2022) produce qualitatively the same results, see the Supplemental Material (SM). As it was previously shown Russ et al. (2020), at large times the generated sequences accurately reproduces many statistical features of the natural sequences used for the training, and have the same probability of being biologically functional in a given experimental platform; hence, the model is generative. Here, we are concerned with what happens at short and intermediate times, where the influence of the ancestral sequence is still important.
@@
III.2 Measures of epistasis
By using the MSA of a given protein family as input data, and inferring the fitness landscape parameters via the DCA model, the authors of Vigué et al. (2022); Bari et al. (2024) have been able to roughly classify the sites of any specific protein sequence belonging to the family in three categories: conserved, mutable, and epistatically constrained. In order to do so, they defined two site-mutability metrics.
@@ III.2 Measures of epistasis
In terms of these quantities, a site $i$ in a given background sequence $A$ can thus be classified as follows:
@@
Furthermore, and most importantly for the present work, Ref. Bari et al. (2024) has considered a given background $A_{0}$ as the ancestral sequence, and starting from it has performed many parallel evolutions in silico, looking at how each site diversifies in the library of mutants obtained after a certain evolutionary time. It was found that mutable sites evolve very rapidly and quickly reach their asymptotic mutability, i.e. the CIE. Conserved sites remain so at any time during evolution, hence do not display interesting dynamics. Epistatically constrained sites, instead, are conserved at short evolutionary times, due to the epistatic constraints in the background of the ancestor, which cause a small CDE. But, as soon as the background sequence mutates enough, they can tolerate more mutations, asymptotically reaching the large CIE that they display in the natural alignment. Their mutation, however, is contingent on several mutations happening in the rest of the sequence, which often require a rather long evolutionary time to take place Bari et al. (2024).
@@
Figure 1: Schematic representation of (a) an MSA in which each site is colored based on its mutability as measured by the $\textrm{CDE}_{i}^{A}$ and (b) the protocol of our evolutionary dynamics with the relevant averages. (c) Depending on the epistatic constraints acting on the ancestral sequence and for equal evolutionary times, we observe a different level of diversity of the evolved sequences (colored), corresponding to a different exploration of the functional sequence space (grey).
III.3 Dynamical fluctuations
@@
Next, we consider a star phylogeny of independent MCMC evolutionary trajectories, all starting from the same ancestor and evolving in parallel (Fig. 1b). Hence, for each ancestor, we construct, as a function of evolutionary time, an MSA of descendants that mimic those obtained through in vitro evolution experiments. As a measure of diversity, we focus on the evolution of the Hamming distance, between the ancestor $A_{0}$ and the MSA of evolving sequences at time $t$, $A_{t}=(a^{t}_{1},\cdots,a^{t}_{L})$, defined as
@@
To make this observation quantitative, we characterize the statistical properties of the resulting MSAs, and how they depend on the ancestor, by introducing two distinct averages and corresponding fluctuations, inspired from the disordered systems literature Franz et al. (2011, 2013); Seoane and Zamponi (2018); Berthier and Jack (2007); Folena et al. (2022) and illustrated in Fig. 1b. Recall that the number of accepted mutations $H^{A_{0}}(t)$ at evolutionary time $t$ for fixed ancestor $A_{0}$ is a random variabile, whose realizations depend on the stochasticity of the evolutionary trajectory.
@@
is the average of the ancestor-dependent $\chi^{A_{0}}_{\rm dyn}(t)$ over $A_{0}$. Using these quantities, which are called ‘dynamical susceptibilities’ in the physics of disordered systems, we can thus quantify the relative importance of the ancestor and the evolutionary noise in determining the diversity of the resulting MSAs at a fixed evolutionary time. In the SM we also study some alternative definitions of susceptibilities, based for example on the Hamming distance between two chains starting from the same ancestor. However, because our MCMC algorithm is time-reversible, these two quantities are related.
IV Results
IV.1 Mutational dynamics and its fluctuations